Calculate the pH after 4.0 mL of 0.50 M NaOH is added to 110.0 mL of a buffer made of 0.40 M CH3COOH and 0.50 M NaCH3COO. The Ka of acetic acid is 1.8 x 10.
Answer #1

first to calculate the moles of each.

then we find moles of NaOH added.

it react with acid.

then to find the moles of acid CH3COOH left.

then to find Molarity in total volume.

then to calculate pH.

Solm Tom Giren, [C13200 1] = 0.4m. Duffes. - [NA Cuzc007 = 0.5M Ku 18x10-5 Jwitiat pt = pta Flog den.com - 5-109[1-8810-5) +lto calulute molarity in total volue = 110 + 4 mL = 0.114L. ch₂coon = 0.04&mol 0 8 0.3684M o-lYL C4₃COO- 0.055 nol = 0.4825 m